How do u do the problem.... find the value of r so that the line throught the pair of points has the given slope. (5,r), (2,-3) with the slope of 4/3
Answer by richwmiller(17219)

(r-(-3))/(5-2)=4/3
(r+3)/3=4/3
r+3=4
r=1